# Contextual AD Narration with Interleaved Multimodal Sequence

### v1

@@ 4.1 Implementation Details
4.1.1 Dataset.
We follow AutoAD-I [15] to conduct partial-data pre-training on the AudioVault-AD dataset [15] and WebVid2M[5]. We train our model on the MAD-v2-Named dataset and evaluate on MAD-eval-Named [40, 15]. More information about dataset is provided in the supplementary material. We use both classic captioning metrics and a recall-based metric for evaluation. The former metrics include ROUGE-L[23] (R-L) and CIDEr[44] (C) to measure the quality of our generated ADs versus human-annotated ones. The latter metric called ‘Recall@$k$ within N neighbours’[14] (R@k/N) calculates the average value of Recall@$k$ for each AD with its $N$ temporally adjacent GT texts. BertScore[48] is used for text similarity measuring.
4.1.2 Training details.
We train our GPT-based model with a batch size of 96 movie clips and the learning rate is $10^{-3}$, while our LLaMA-based model is trained with a batch size of 12 movie clips and the learning rate is $5.0\times 10^{-5}$. We use the AdamW[27] optimizer to train our model for 10 epochs, with a cosine-decayed learning rate schedule and linear warm-up. All training is conducted on 8 A100-80G GPUs.
@@ 4.2 Comparison with state-of-the-art approaches
Table 1: Comparison with the state-of-the-art methods on MAD-eval-Named. The Context column denotes whether contextual information is applied. The V-Feature column indicates the type of visual expert used for extracting movie frame features. AV denotes AudioVault-AD and WV denotes WebVid2M.
Methods Context Pre-train LLM & V-Feature RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
ClipCap[28] CC3M GPT-2 & CLIP-B32 8.5 4.4 36.5
AutoAD-I[15] AV & WV GPT-2 & CLIP-B32 10.3 12.1 39.8
AutoAD-II[14] AV & WV GPT-2 & CLIP-B32 13.1 19.2 51.3
Uni-AD(ours) AV & WV GPT-2 & CLIP-B32 15.9 24.0 50.5
Uni-AD(ours) AV & WV GPT-2 & CLIP-L14 16.4 25.7 51.5
Uni-AD(ours) AV LLaMA-7B & CLIP-L14 16.8 27.3 53.3
AutoAD-I[15] AV & WV GPT-2 & CLIP-B32 11.9 14.3 42.1
AutoAD-II[14] AV & WV GPT-2 & CLIP-B32 13.4 19.5 50.8
MM-Narrator[46] - GPT-4 & CLIP-L14 13.4 13.9 49.0
MM-Narrator[46] - GPT-4V & CLIP-L14 12.8 9.8 -
Uni-AD(ours) AV LLaMA-7B & CLIP-L14 17.2 28.2 54.9
Multiple models for AD generation are involved in our comparison. Descriptions for these methods are available in the supplementary material. Tab. 1 shows the evaluation results of our Uni-AD and these state-of-the-art methods on the MAD-eval-Named benchmark. For fair comparison, we first evaluate Uni-AD under the same setting (GPT-2 as language decoder and CLIP ViT-B/32 as visual feature) with previous methods[28, 15, 14]. Our Uni-AD outperforms all previous methods by a notable margin, which demonstrates the effectiveness of our model. Specifically, we surpass the previously best method AutoAD-II[14] by 2.8 and 4.8 points on metrics Rouge-L and Cider, respectively. Afterwards, we apply a stronger CLIP ViT-L/14 model as visual encoder to extract movie features, and the results of Uni-AD get growth on all metrics. Utilizing the design of lightweight visual mapping network, we then extend our method to a more powerful LLM LLaMA2-7B, which further enhance the performance (1.6 growth points on Cider, 0.4 on Rouge-L and 1.8 on Recall@5/16).
Next, we conduct comparison with AD generation approaches under the contextual setting. Our Uni-AD achieves the state-of-the-art performance (17.2 on Rouge-L, 28.2 on Cider and 54.9 on Recall@5/16) by incorporating context video features and contrastive leaning. This outstanding performance shows the effectiveness and flexibility of our model.
@@ 4.3 Ablation Study
Table 2: Study the effectiveness of character-refinement module without pre-training or contextual information. Char.? shows whether this module is applied. gt indicates directly using characters in AD annotation as character information. * denotes our re-implemented results.
Methods Char.? RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
AutoAD-II[14] (GPT-2-B32)    $*{}^{*}$    $*{}^{*}$    $*{}^{*}$
@@
Methods #Latent RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@
Table 2: Study the effectiveness of character-refinement module without pre-training or contextual information. Char.? shows whether this module is applied. gt indicates directly using characters in AD annotation as character information. * denotes our re-implemented results.
Table 3: Study the impact of visual mapping network on AD generation without pre-training or contextual information. character-refinement module is applied by all methods. The #Latent column denotes the number of learnable vectors in the visual mapping network.
4.3.1 Study on character-refinement module.
By recognizing the main characters contributed to the storyline, the character-refinement module provides more precise character information for AD generation. Evaluation results in Tab. 3 show the general performance improvements on all metrics character-refinement module brings under different settings, confirming the universality of this module. Note that when character-refinement module is not applied, the Flamingo-based method AutoAD-II achieves better results (e.g. 14.7 versus 14.0 on Rouge-L), while with refined character information, our Uni-AD surpasses AutoAD-II (15.7 versus 15.1 on RL, 23.7 versus 21.0 on C, 49.4 versus 47.8 on R@5/16). We speculate that this is because AutoAD-II incorporates learnable parameters at multiple locations within GPT-2, thus allowing for continuous adjustment of the generated output to reduce the impact of noise. When the input is more precise, our model can make fuller use of this information to produce better results.
We further use characters involved in AD annotations as character information to conduct experiments (gt rows in Tab. 3). Once again, Uni-AD achieves the best performance. The gap between character-refinement results and gt results also show that the current character-refinement performance (0.41 on Precision, 0.77 on Recall, evaluated on MAD-eval-Named dataset) is far from sufficient and there is still significant room for improvement.
4.3.2 Study on Visual Mapping Network.
To verify whether our Uni-AD can achieve finer-grained feature alignment, we in this section study how the visual mapping network affects AD generation. Our main comparison target is the Flamingo-style method AutoAD-II. Specifically, we adjust the number of latent vectors output by the visual mapping network, which also corresponds to the number of visual tokens fed to the LLM, to observe the impact on AD generation. The evaluation results are shown in Tab. 3. It can be seen that as the number of latent vectors increases, the performance of Uni-AD improves under both settings (GPT-CLIP-B32 and LLaMA-CLIP-L14), while the result of AutoAD-II method remains basically unchanged. This confirms our hypothesis that AutoAD-II, which achieves cross-modal alignment by concatenating characters’ portraits with video frames followed by perceiver resampler and gated cross-attention modules, tends to extract global feature of the visual contents. In contrast, our Uni-AD is capable of retaining more visual details by increasing the number of latent vectors, thereby achieving finer-grained feature alignment and better AD generation results.
4.3.3 Study on Partial-Data Pre-training.
Methods Pre-train RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
AutoAD-II [14] None 15.1 21.0 47.8
AudioVault-AD & WebVid 15.2 21.7 49.3
Uni-AD (GPT-2-B32) None 15.7 23.7 49.4
AudioVault-AD & WebVid 15.9 24.0 50.5
Uni-AD (GPT-2-L14) None 16.0 25.5 50.3
AudioVault-AD & WebVid 16.4 25.7 51.5
Uni-AD (LLaMA-L14) None 16.5 25.9 52.5
AudioVault-AD 16.8 27.3 53.3
WebVid 12.6 15.9 39.4
AudioVault-AD & WebVid 12.4 16.6 39.8
Table 4: Study the impact of partial-data pre-training on AD generation. All results are obtained with the character-refinement module.
To address the issue of insufficient paired video-AD data, AutoAD-I[15] proposed to fine-tune GPT-2 on text-only AudioVault-AD[15] dataset and pre-train visual mapping module on WebVid2M[5]. In this section, we follow this idea to perform partial data pre-training for our method. Note that to reduce memory usage, we apply LoRA[16] to fine-tune our LLaMA-CLIP-L14 model on AudioVault-AD. Tab. 4 shows that all GPT-2 based models can be further improved by this method, while pre-training on WebVid leads to performance-degradation in our LLaMA-based model. We attribute this to the large domain and quantity gap between MAD and WebVid. Given the large number of parameters in LLaMA, it is hard to align the visual module with LLaMA in a new feature space using data from MAD after pre-training on the large-scale dataset WebVid.
@@
(a) Study the impact of context video and contrastive loss. Context-V denotes the number of past context video clips. C-Loss indicates whether contrastive loss is used.
Context-V C-Loss RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@
(b) Study the effectiveness of context AD. Context-AD denotes the number of past context ADs. V&\&&C-Loss indicates whether past videos and contrastive loss are applied.
Context-AD V&C-Loss RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@
4.4 Integrating Contextual Information
For contextual information, we first study the effectiveness of integrating context video and contrastive loss into AD generation. Results in Tab. 5(a) suggest that adding context video clips and contrastive loss can both enhance our model’s performance. Among these results, incorporating the most recent context video along with contrastive loss is the best, since current AD is always most relevant to its preceding content.
Then we conduct experiments to introduce past ADs as additional contextual information under recurrent setting (using predicted past ADs) and oracle setting (using ground-truth past ADs). As shown in Tab. 5(b), the recurrent setting leads to a decrease in model performance. We attribute this to the discrepancy between the predicted ADs and the ground truth, which provides inaccurate information for AD generation during inference, thus further enlarges the gap between training and testing. For oracle setting, the performance of our model improves with more context ADs, indicating that context AD is the most significant influencing factor for AD generation.
@@ 5 Conclusion
In this work, we present a simple and unified framework called Uni-AD for Audio Description (AD) generation task by prompting pre-trained LLMs with interleaved multimodal sequence as input. Compared with previous work, our Uni-AD is able to leverage more precise character information provided by the character-refinement module and fully utilize rich contextual information to generated ADs. Uni-AD achieves the state-of-the-art performance on the AD generation benchmark MAD-named. We also conduct comprehensive ablation studies to validate the effectiveness of different components, which demonstrates that fine-grained feature alignment, precise character information, and contextual data can benefit AD generation. We hope our work could facilitate research in this community.
@@
The appendix provides more implementation details, ablation analyses and qualitative results to show deep insights into our method. Specifically,  Sec. A describes the architecture details of our model and experiment settings.  Section B provides more ablation studies on our Uni-AD. To further show the effectiveness of our model design, we show more qualitative results of Uni-AD in  Sec. C.
A Implementation Details
A.1 Architecture details
We show the architecture details of visual mapping network and character-refinement module in  Tab. 6.
@@ A.1 Architecture details
Table 6: Architecture details of visual mapping network and character-refinement module in Uni-AD.
@@
Visual mapping network (GPT-2-B32) Visual mapping network (LLaMA-L14) Character-refinement module
projection layer 512$\rightarrow$768 768$\rightarrow$4096 -
num latent 30 30 -
num blocks 2 2 3
channel 768 4096 768
num head 12 32 12
ffn dimension 3072 16384 3072
A.2 Datasets
A.2.1 MAD-Named
@@
[40, 15]: The MAD-Named benchmark consists of two parts: MAD-v2-Named for training and MAD-eval-Named for testing. Specifically, MAD-v2-Named contains 334,296 ADs and 628,613 subtitles collected from 488 movies, while MAD-eval-Named cotains 6,520 ADs and 10,602 subtitles collected from 10 movies. Annotation include the start and end time of each AD and the AD contents without any post-processing on the character names. We notice that there are many homophonic name mismatches in MAD-Named, such as an actor’s name being ’Gray’ in the character information from IMDb but ’Grey’ in the ad annotation. Thus we have processed these mismatched information to ensure that the same actor’s name remains consistent.
A.2.2 AudioVault-AD
@@
[15]: AudioVault-AD is a text-only dataset composed of 3.3 million AD utterances collected from 7,057 movies downloaded from the AudioVault website. Movies in AudioVault-AD are not included in the MAD dataset.
A.2.3 WebVid2M
[5]: WebVid2M is a large video-caption dataset that contains 2.5M short stock footage videos, which we applied for partial-pretraining of visual module.
A.3 Baselines
In this section, we compare our Uni-AD with recent AD generation methods:
A.3.1 ClipCap
@@
[28]. The ClipCap model converts the CLIP[35] feature of visual inputs into embeddings with a mapping network. Then the output embeddings will be used as prefix to prompt GPT-2[36] to generate corresponding captions.
A.3.2 AutoAD-I
@@
[15]. AutoAD-I follows ClipCap and concatenate previous AD descriptions and subtitles in movie with visual embeddings to prompt the fronzen GPT-2[36] for AD generation. This approach further apply partial-data pretrain to address the issue of insufficient AD data.
A.3.3 AutoAD-II
@@
[14]. This method applies a Flamingo-style[4] architecture for AD generation and introduces an external Character Bank to enable their model to label characters appearing in the movie. AutoAD-II also presents an AD temporal proposal module to determine whether AD should be inserted in the given pause in dialogue.
A.3.4 MM-Narrater
@@
[46]. MM-Narrater employs specialized vision and audio expert models to extract multimodal information from the input video clip. The outputs along with movie subtitles and previous AD descriptions are used to build prompt to query GPT-4[31] or GPT-4V[32] for AD generation. Besides, MM-Narrater utilizes retrieval enhancement and in context learning to improve the quality of generated AD.
B More Ablation Studies

### v2

@@ 4.1 Implementation Details
Dataset. We follow AutoAD-I [15] to conduct partial-data pre-training on the AudioVault-AD dataset [15]. We train our model on the MAD-v2-Named dataset and evaluate on MAD-eval-Named [41, 15]. For evaluation on CMDAD [17] and TVAD [48], we train our model with the CMDAD [17] training set. We use both classic captioning metrics and newly proposed metrics for evaluation. The former metrics include ROUGE-L [24] (R-L) and CIDEr [45] (C) to measure the quality of our generated ADs versus human-annotated ones. The latter metrics include R@k/N [16], CRITIC [17] and LLM-AD-eval [17]. More information about datasets and metrics is provided in the supplementary material.
Training details. We train our GPT-based model with a batch size of 96 movie clips and the learning rate is $10^{-3}$, while our LLaMA-based model is trained with a batch size of 12 movie clips and the learning rate is $5.0\times 10^{-5}$. We use the AdamW [28] optimizer to train our model for 10 epochs, with a cosine-decayed learning rate schedule and linear warm-up. All training is conducted on 8 A100 GPUs. For external character information, we use the prediction results from AutoAD-II [16] as input for our character refine module on MAD-eval-Named dataset and prediction results from AutoAD-Zero [48] as input for character refine module on CMDAD and TVAD. We apply the visual branch of VideoLLaMA [50] as visual mapping network on experiments of CMDAD and TVAD, as in  [17].
@@ 4.2 Comparison with state-of-the-art approaches
Table 1: Comparison with the state-of-the-art methods on MAD-eval-Named. The Context column denotes whether contextual information is applied. The V-Feature column indicates the type of visual expert used for extracting movie frame features.
[…]
ClipCap [29] GPT-2 & CLIP-B32 8.5 4.4 36.5
AutoAD-I [15] GPT-2 & CLIP-B32 10.3 12.1 39.8
AutoAD-II [16] GPT-2 & CLIP-B32 13.1 19.2 51.3
AutoAD-III [17] LLaMA2 & EVA-CLIP - 24.0 52.8
AutoAD-Zero [48] LLaMA3 & VideoLLaMA2 - 22.4 -
Uni-AD(ours) GPT-2 & CLIP-B32 15.9 24.0 50.5
Uni-AD(ours) GPT-2 & CLIP-L14 16.4 25.7 51.5
Uni-AD(ours) LLaMA2 & CLIP-L14 16.8 27.3 53.3
AutoAD-I [15] GPT-2 & CLIP-B32 11.9 14.3 42.1
AutoAD-II [16] GPT-2 & CLIP-B32 13.4 19.5 50.8
MM-Narrator [49] GPT-4 & CLIP-L14 13.4 13.9 49.0
MM-Narrator [49] GPT-4V & CLIP-L14 12.8 9.8 -
Uni-AD(ours) LLaMA2 & CLIP-L14 17.1 28.2 54.2
Table 2: Comparison with the state-of-the-art methods on CMDAD and TVAD. The gray row shows the results of AutoAD-III pretrained on the 3.4M HowTo-AD dataset, which is not public.
[…]
AutoAD-II [16] CMDAD 13.5 8.2 2.08
AutoAD-III [17] CMDAD 21.7 25.2 2.85
AutoAD-III [17] CMDAD 25.0 32.7 2.92
AutoAD-Zero [48] CMDAD 17.7 43.7 2.83
Uni-AD(Ours) CMDAD 21.8 41.9 2.92
AutoAD-III [17] TVAD 26.1 15.9 2.78
AutoAD-Zero [48] TVAD 22.6 27.6 2.94
Uni-AD(Ours) TVAD 26.6 28.3 2.89
Multiple models for AD generation are involved in our comparison. Descriptions for these methods are available in the supplementary material. Tab. 1 shows the evaluation results of our Uni-AD and these state-of-the-art methods on the MAD-eval-Named benchmark. For fair comparison, we first evaluate Uni-AD under the same setting (GPT-2 as language decoder and CLIP ViT-B/32 as visual feature) with previous methods [29, 15, 16]. Our Uni-AD outperforms all previous methods by a notable margin, which demonstrates the effectiveness of our model. Afterward, we apply a stronger CLIP ViT-L/14 model as visual encoder to extract movie features, and the results of Uni-AD show growth on all metrics. Utilizing the design of lightweight visual mapping network, we then extend our method to a more powerful LLM LLaMA2-7B, which further enhances the performance (1.6 growth points on Cider, 0.4 on Rouge-L and 1.8 on Recall@5/16).
Next, we conduct comparison with AD generation approaches under the contextual setting. Uni-AD achieves the state-of-the-art performance (17.1 on Rouge-L, 28.2 on Cider and 54.2 on Recall@5/16) by incorporating context video features and contrastive learning. We further evaluate Uni-AD on CMDAD and TVAD and obtain competitive results, shown in Tab. 2. These outstanding performances show the effectiveness and flexibility of our model.
@@ 4.3 Ablation Study
Table 3: Study the effectiveness of character-refinement module without pre-training or contextual information. Char.? shows whether this module is applied. * means our re-implemention.
[…]
AutoAD-II (GPT-2-B32) [16]    14.7    19.0    46.0
@@
Study on character-refinement module. By recognizing the main characters contributed to the storyline, the character-refinement module provides more precise character information for AD generation. Evaluation results in Tab. S2 show the general performance improvements character-refinement module brings under different settings, confirming the universality of this module.
We further use characters involved in AD annotations as character information to conduct experiments (gt rows in Tab. S2). The gap between character-refinement results and gt results also show that the current character-refinement performance (0.41 on Precision, 0.77 on Recall, evaluated on MAD-eval-Named dataset) is far from sufficient and there is still significant room for improvement.
Study on Visual Mapping Network. To verify whether our Uni-AD can achieve finer-grained feature alignment, we in this section study how the visual mapping network affects AD generation. Our main comparison target is the Flamingo-style method AutoAD-II. Specifically, we adjust the number of latent vectors output by the visual mapping network, which also corresponds to the number of visual tokens fed to the LLM, to observe the impact on AD generation. The evaluation results are shown in Tab. 4. It can be seen that as the number of latent vectors increases, the performance of Uni-AD improves under both settings (GPT-CLIP-B32 and LLaMA-CLIP-L14), while the result of AutoAD-II method remains basically unchanged. This confirms our hypothesis that AutoAD-II, which achieves cross-modal alignment by concatenating characters’ portraits with video frames followed by perceiver resampler and gated cross-attention modules, tends to extract global feature of the visual contents. In contrast, our Uni-AD is capable of retaining more visual details by increasing the number of latent vectors, thereby achieving finer-grained feature alignment and better AD generation results.
Table 4: Study the impact of visual mapping network on AD generation without pre-training or Contextual information. Character-refinement module is applied by all methods. #Latent denotes the number of learnable vectors in the visual mapping network.
[…]
@@
[…]
@@
4.4 Integrating Contextual Information
For contextual information, we first study the effectiveness of integrating context video and contrastive loss into AD generation. Results in Tab. 5 suggest that adding context video clips and contrastive loss can both enhance model’s performance. Among these results, incorporating the most recent context video along with contrastive loss is the best, since current AD is most relevant to its preceding content.
Then we conduct experiments to introduce past ADs as additional contextual information under recurrent setting (using predicted past ADs) and oracle setting (using ground-truth past ADs). As shown in Tab. 6, the recurrent setting leads to a decrease in model performance. We attribute this to the discrepancy between the predicted ADs and the ground truth, which provides inaccurate information for AD generation during inference, thus further enlarges the gap between training and testing. For oracle setting, the performance of our model improves with more context ADs, indicating that context AD is the most significant influencing factor for AD generation.
Table 6: Study the effectiveness of context AD. Context-AD denotes the number of past context ADs. V&\&&C-Loss indicates whether past videos and contrastive loss are applied.
[…]
@@ 5 Conclusion
In this work, we present a simple and unified framework called Uni-AD for Audio Description (AD) generation task by prompting pre-trained LLMs with interleaved multimodal sequence as input. Compared with previous work, our Uni-AD is able to leverage more precise character information provided by the character-refinement module and fully utilize rich contextual information to generated ADs. Uni-AD achieves the state-of-the-art performance on multiple AD generation benchmarks. We also conduct comprehensive ablation studies to validate the effectiveness of different components, which demonstrates that fine-grained feature alignment, precise character information, and contextual data can benefit AD generation. We hope our work could facilitate research in this community.
Acknowledgements: This work is supported by the National Key R&\&&D Program of China (No. 2022ZD0160900), ERC AdG Project (No. 101021347), Ant Group Research Intern Program, Jiangsu Frontier Technology Research and Development Program (No. BF2024076), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.
[…]
Supplementary Material
@@ Appendix A Overview
The supplementary material provides more implementation details, ablation analyses and qualitative results to show deep insights into our method. Specifically,  Appendix B describes the architecture details of our model and experiment settings.  Appendix C provides more ablation studies on our Uni-AD. To further show the effectiveness of our model design, we show more qualitative results of Uni-AD in  Appendix D.
Appendix B Implementation Details
B.1 Architecture details
We show the architecture details of our visual mapping network for MAD-eval-Named benchmark and character-refinement module in  Tab. S1. Note that for CMDAD and TVAD datasets, we use the training set of CMDAD to train our character-refinement module and apply InternVideo2.5 [47] to extract video frame features. The character-refinement performances are shown in  Tab. S2.
@@ B.1 Architecture details
Table S1: Architecture details of visual mapping network and character-refinement module in Uni-AD.
Visual mapping network (GPT-2-B32) projection layer 512$\rightarrow$768
num latent 30
num blocks 2
channel 768
num head 12
ffn dimension 3072
Visual mapping network (LLaMA-L14) projection layer 768$\rightarrow$4096
num latent 30
num blocks 2
channel 4096
num head 32
ffn dimension 16384
Character-refinement module (for MAD-eval-Named) num blocks 3
channel 768
num head 12
ffn dimension 3072
Character-refinement module (for CMDAD&TVAD) projection layer 4096$\rightarrow$4096
num blocks 1
channel 4096
num head 32
ffn dimension 16384
@@
Table S2: Architecture details of visual mapping network and character-refinement module in Uni-AD.
[…]
MAD-eval-Named 0.41 0.77
CMDAD 0.26 0.94
TVAD 0.27 0.94
B.2 Datasets
@@
MAD-Named [41, 15]: The MAD-Named benchmark consists of two parts: MAD-v2-Named for training and MAD-eval-Named for testing. Specifically, MAD-v2-Named contains 334,296 ADs and 628,613 subtitles collected from 488 movies, while MAD-eval-Named contains 6,520 ADs and 10,602 subtitles collected from 10 movies. Annotation includes the start and end time of each AD and the AD contents without any post-processing on character names. We notice there are many homophonic name mismatches in MAD-Named, such as an actor’s name being ’Gray’ in character information from IMDb but ’Grey’ in the ad annotation. We thus processed these mismatched information to ensure that the same actor’s name remains consistent.
CMDAD [17]: CMDAD is a movie AD dataset that contains 101k ADs for more than 1432 movies, with 100 movies split for evaluation.
TVAD [48]: TVAD is a recently proposed TV-series AD dataset, which contains 31k ADs for training and 3k ADs for evaluation.
@@
AudioVault-AD [15]: AudioVault-AD is a text-only dataset composed of 3.3 million AD utterances collected from 7,057 movies downloaded from the AudioVault website. Movies in AudioVault-AD are not included in the MAD dataset.
B.3 Baselines
In this paper, we compare our Uni-AD with the following AD generation methods:
@@ B.3 Baselines
ClipCap [29]. The ClipCap model converts the CLIP [37] feature of visual inputs into embeddings with a mapping network. Then the output embeddings will be used as prefix to prompt GPT-2 [36] to generate corresponding captions.
@@
AutoAD-I [15]. AutoAD-I follows ClipCap and concatenate previous AD descriptions and subtitles in movie with visual embeddings to prompt the fronzen GPT-2 [36] for AD generation. This approach further apply partial-data pretrain to address the issue of insufficient AD data.
@@
AutoAD-II [16]. This method applies a Flamingo-style [4] architecture for AD generation and introduces an external Character Bank to enable their model to label characters appearing in the movie. AutoAD-II also presents an AD temporal proposal module to determine whether AD should be inserted in the given pause in dialogue.
AutoAD-III [17]. AutoAD-III follows BLIP2 [22] to use Q-former architecture to bridge the visual space with the language space. Then the model can generate textual outputs with a large language model. AutoAD-III also proposed a large-scale HowTo-AD dataset for pre-training.
AutoAD-Zero [48]. AutoAD-Zero designs a pipeline for character recognition with face detection methods and prompts LLM by circling character faces. A two-stage training-free method is proposed for AD generation, which consists of (i) VLM-Based Video Description and (ii) LLM-Based AD Summary.
@@
MM-Narrater [49]. MM-Narrater employs specialized vision and audio expert models to extract multimodal information from the input video clip. The outputs, along with movie subtitles and previous AD descriptions are used to build prompt to query GPT-4 [32] or GPT-4V [33] for AD generation. Besides, MM-Narrater utilizes retrieval enhancement and in context learning to improve the quality of generated AD.
B.4 Metircs
In this paper, we use both classic captioning metrics and newly proposed metrics for evaluation. Classic captioning metrics include ROUGE-L [24] and CIDEr [45]. In this section, we mainly introduce newly proposed metrics: R@k/N, CRITIC and LLM-AD-eval.
R@k/N [16]: R@k/N is a retrieval metric that distinguishes the predicted text among a set of neighbours. The parameters $k$ and $N$ mean within a temporal window of $N$ neighbouring reference ADs, whether the predicted AD can retrieve the corresponding reference AD at top-k position.
CRITIC [17]: CRITIC assesses the precision of character recognition in generated ADs. Specifically, a co-referencing model is utilized to substitute ambiguous pronouns in ADs with official names from the character banks. Subsequently, two sets of names from predicted and ground truth ADs are compared, and the IoU is computed to yield a CRITIC score.
LLM-AD-eval [17]: LLM-AD-eval utilises LLMs to judge the quality of generated ADs by scoring them between 1 (lowest) and 5 (highest). We use llama2-7b-chat [44] for the evaluation in our experiments.
Appendix C More Ablation Studies